OpenStudy (anonymous):

How many ways can you get a sum of 7 or more from a roll of two dice?

OpenStudy (anonymous):

Since there are only \(6\times 6=36\) combinations total, it wouldn't be hard to count which add up to 7.

OpenStudy (anonymous):

Yeah, 16 ways, starting with (4,3), (4,4), (4,5).....(5,6), (6,6).

OpenStudy (anonymous):

But i need to know how to display it algebraically

OpenStudy (anonymous):

18 ways, my apologies

OpenStudy (anonymous):

Okay, well consider what the first die is, and how many ways you can get a sum > 6 in each case. With 6 there are six ways, 5 there are 5, 4, there are 4, etc. So it's just: \[\large \sum_{k=1}^nk^2 = \frac{1}{6}n(n+1)(2n+1)\]In our case, \(n=6\).

OpenStudy (anonymous):

Wow, so that is the binomial type function that represents t. Thanks you very much.

OpenStudy (anonymous):

Is there some other combinatoric rule you are supposed to use?

OpenStudy (anonymous):

I don't believe so. We are just working with permutations, combinations and factorials